OpenStudy (anonymous):

Very first problem set. How do I get the program to print out the birthdate and last name together? Right now I have: print (raw_input('enter your date of birth')) print (raw_input('enter your last name'))

OpenStudy (anonymous):

Have you read any of the readings?

OpenStudy (anonymous):

the wikilinks? I read over those, but am brand new to any and all of this, so I've been struggling to get started.

OpenStudy (anonymous):

Which class are you doing? 2008?, 2011?

OpenStudy (anonymous):

the 2011 course

OpenStudy (anonymous):

Have you read the Tutorial in the python docs? Tried all of the examples?

OpenStudy (anonymous):

here are a couple more but maybe start with the Tutorial in the documentation http://troll.cs.ua.edu/cs150/book/index.html http://www.greenteapress.com/thinkpython/html/index.html you are going to need to get used to the documentation and start finding answers. when you want to reuse the result of a statement, you 'store' it in a variable ``` >>> >>> a = 2 + 3 >>> b = 2 - 3 >>> print a, b 5 -1 >>> ``` a and b are variables

OpenStudy (anonymous):

okay thank you. i'll keep working on it

OpenStudy (anonymous):

I did not need to put the print command on the first two strings in order to get it to function

OpenStudy (anonymous):

nope ... it helps if your variable have descriptive names for short pieces of code or code examples, you can enclose it in 'bacticks' (left of 1 on a qwerty keyboard.) backticks code backticks ``` >>> dob = (raw_input('enter your date of birth')) enter your date of birth5/5/5555 >>> last = raw_input('enter your last name') enter your last namefinn >>> first = raw_input('enter your first name') enter your first namealva >>> print first, last, dob alva finn 5/5/5555 >>> ``` Otherwise please use a code pasting site: - http://dpaste.com - http://pastebin.com - https://gist.github.com/ - http://pastie.org - http://codepad.org - http://ideone.com - http://www.repl.it/ paste your code there and post the link here. select Python syntax highlighting when u paste. ... http://dpaste.com/1406703/